# Defending against adversarial attacks using mixture of experts

### v1

@@ Abstract
Machine learning is a powerful tool enabling full automation of a huge number of tasks without explicit programming. Despite recent progress of machine learning in different domains, these models have shown vulnerabilities when they are exposed to adversarial threats. Adversarial threats aim to hinder the machine learning models from satisfying their objectives. They can create adversarial perturbations, which are imperceptible to humans’ eyes but have the ability to cause misclassification during inference. Moreover, they can poison the training data to harm the model’s performance or they can query the model to steal its sensitive information. In this paper, we propose a defense system, which devises an adversarial training module within mixture-of-experts architecture to enhance its robustness against adversarial threats. In our proposed defense system, we use nine pre-trained experts with ResNet-18 as their backbone. During end-to-end training, the parameters of expert models and gating mechanism are jointly updated allowing further optimization of the experts. Our proposed defense system outperforms state-of-the-art defense systems and plain classifiers, which use a more complex architecture than our model’s backbone.
@@ 1 Introduction
Machine learning (ML) is a cornerstone of artificial intelligence (AI), which has been widely used in numerous applications. In computer vision, it enables face recognition and video surveillance [khan2021machine, khan2020machine]. In health care, it facilitates disease diagnosis, drug repurposing, and personalized treatments [habehh2021machine, qayyum2020secure]. In natural language processing (NLP), ML enhances the quality of language translation [lauriola2022introduction, omar2022robust]. ML also empowers Large Language Models (LLMs) for code-related tasks such as code generation and understanding [jelodar2025large]. In cybersecurity, ML eases threat detection and manages system’s vulnerabilities [apruzzese2023role, xin2018machine, buczak2015survey]. Moreover, ML also plays an important role in other domains such as finance, speech processing, recommendation systems, autonomous vehicles, among others. However, those wide range of applications of machine learning has introduced opportunities for adversaries to exploit natural vulnerabilities of these models and launch adversarial attacks to threaten the integrity, privacy, and availability of these models [pitropakis2019taxonomy, bountakas2023defense, kurakin2016adversarial, kaur2022trustworthy].
Adversarial machine learning (AML) is at intersection of cybersecurity and machine learning and aims to investigate attacks on machine learning models, and design defense systems to counter these attacks. The adversary may have some knowledge about the victim’s model and/or its training data, which facilitates the adversarial sample generation [pitropakis2019taxonomy, bountakas2023defense, kurakin2016adversarial]. Figure 1 shows how an adversarial sample looks like on an MNIST [lecun2002gradient] sample when the adversary has a perfect knowledge about the model’s parameters. Samples like this can cause model’s misclassification with a high confidence during test time.
@@
The rivalry between governments and industrial corporations has given rise to adversarial attacks causing global concerns about adversarial machine learning. For instance, the government of United States of America has established Bipartisan AI task force, which are responsible for appropriate use of AI. They introduced AI guidelines such as accuracy, reliability, robustness, privacy, transparency, and safety. They also labeled generative and adversarial AI as a threat to privacy and confidentiality [AI_Task_Force_2024]. Moreover, the Department of Homeland Security (DHS) has considered evasion attacks and generative AI as a national threat violating security and privacy of individuals [DHS_AI_Threat_2023]. These insights highlight the need for secure AI systems and robust ML models.
Trustworthy AI is an important domain of AI ensuring that AI models remain lawful, ethical, robust, and efficient. Robust machine learning is an important branch in trustworthy AI which aims to make the machine learning models robust against adversarial threats while remain functional in non-adversarial environments [kaur2022trustworthy]. Moreover, the main focus of Robust ML is to develop defense systems that detect and/or mitigate adversarial threats while preserving their efficiency and performance to a certain extent. These defense systems could either reside in the models or serve as an external entity protecting ML models [pitropakis2019taxonomy, bountakas2023defense, kurakin2016adversarial, kaur2022trustworthy].
In pursuit of building a secure system, we build a novel defense system named Divided We Fall (DWF). In DWF, we employ Mixture of Experts (MoE) architecture which consists of benign and adversarial experts. The training scheme in DWF is joint, meaning that we do not freeze the weights of our pretrained experts during end-to-end training, which leads to stronger generalization capability of our proposed defense system improving the clean accuracy and robustness of our model.
@@ 2 Background and Problem Setting
MoE is an ensemble architecture using multiple experts, a gating mechanism to assign weights, and a router to aggregate results [mu2025comprehensive, cai2025survey, puigcerver2023sparse]. MoEs can be categorized based on their gating mechanism, routing topology, training strategy, and experts activation, which is shown in Figure 2.
Based on gating mechanism, MoE could be either soft or hard; in soft MoE the gating assigns non-zero weights to each expert forming fractional activation of experts, while in hard MoE, the gating mechanism chooses a subset of experts and assign wights to them making experts activation as sparse, or single if only one of them is chosen [mu2025comprehensive, cai2025survey, puigcerver2023sparse].
Training strategy could be either joint or frozen. In the first one, both router and experts are trained simultaneously, while the latter one uses pretrained experts and freezes their weights during training of the router [mu2025comprehensive, cai2025survey, puigcerver2023sparse].
Based on the routing topology we have flat and hierarchical routing schemes. In the flat MoE, all the experts reside in a single layer and the router chooses from them. In the hierarchical topology, the first router groups the experts into subgroups, and, then, the second one chooses some experts from these subgroups [mu2025comprehensive, cai2025survey].
@@
Adversarial threats can be viewed from different angles: attacker’s knowledge, type, specificity, and objective. Figure 3 shows an overview of adversarial threat model.
2.1 Attacker’s Knowledge
Attacker’s knowledge shows the access level of the adversary to model’s parameters, gradients, architecture, and training data. This knowledge is categorized into white-box, black-box, and gray-box based on the adversary’s access level [pitropakis2019taxonomy, bountakas2023defense].
2.1.1 White Box
Adversaries with white-box knowledge are the most effective and destructive attackers. These attackers have a full access to model’s architecture, parameters, and gradients during training with full knowledge on the training data; this enables the attackers to generate adversarial samples based on the gradients and training data, which makes them very powerful. However, in a real-world scenario, adversaries can’t achieve such an access, making them more suitable for research and testing the models’ robustness [pitropakis2019taxonomy, bountakas2023defense].
2.1.2 Black Box
In black-box scenario, unlike white-box, the adversaries have no knowledge about the model and the presence of any defense system. The adversaries can only give inputs and receive the corresponding outputs [pitropakis2019taxonomy, bountakas2023defense, lin2023sensitive].
2.1.3 Gray Box
Gray box knowledge lies between white box knowledge and black box knowledge. The adversary has some degree of knowledge about the target system, which is not as perfect as white box, and is not as scarce as black box knowledge [pitropakis2019taxonomy, bountakas2023defense].
2.2 Attacker’s specificity
Based on specificity, each attack could either be targeted or untargeted. In targeted attacks, the adversary aims to degrade the model’s performance when it faces specific patterns/inputs, while in untargeted attacks, also named to as indiscriminate, the goal is to cause any misclassification [pitropakis2019taxonomy, bountakas2023defense].
2.3 Attacker’s objective
Based on the objective, the adversary can follow degrade, breach, or both scenarios. In the first scenario, the adversary aims to harm the system’s performance and/or violets its integrity. In breach, the adversary aims to explore the model in order to steal sensitive information, without affecting the model’s performance [pitropakis2019taxonomy, bountakas2023defense].
2.4 Attacker’s type
@@
2.4.1 Evasion
An evasion attack targets the model at test time without affecting the system’s integrity on clean data. Evasion attacks are the most practical attack types aiming to degrade the model’s performance by causing misclassification during testing [pitropakis2019taxonomy, bountakas2023defense, gong2024improving]. For instance attacks such as Fast Gradient Sign Method (FGSM) [goodfellow2014explaining], Projected Gradient Descent (PGD) [madry2017towards], and Carlini & Wagner (C&W) [carlini2017towards] make use of the model’s gradient during training to generate samples which evade the classifier during the test time. Equations 2 and 3, in Section 5, show how an adversarial sample is generated in FGSM and PGD, respectively.
2.4.2 Poisoning
Poisoning attacks aim to violate the system’s integrity by corrupting the training data and/or model embeddings to harm the model integrity. These attacks could be either label poisoning attacks or clean label attacks, where only the former manipulates the labels. Additionally, backdoor poisoning attacks insert hidden triggers/patterns in the training data forcing the model to cause misclassification when some target patterns are observed by the model without affecting model’s overall performance prominently [pitropakis2019taxonomy, bountakas2023defense, tian2022comprehensive, bai2024backdoor, zhang2025apbam, zheng2022concealed].
2.4.3 Exploratory
Exploratory attacks aim to violate privacy of the system. These attacks probe the target model to gain information about model’s parameters, training data, and so on without affecting the performance and integrity of the model [rigaki2023survey, zhang2024defending]. One of the most well-known attacks from this category is named Membership Inference Attack (MIA) [shokri2017membership], which aims to assess the information leakage of a machine learning model by giving an input and receiving the corresponding output in a black-box manner.
2.4.4 Physical
These attacks are launched on the models, which are deployed on a special physical hardware such as FPGA, GPU, edge devices and so on. Attacks on physical layer exploit vulnerabilities, which are emerged due to the increased attack surface in the physical layer [regazzoni2020machine]. For instance, in a power analysis attack, the adversary can infer cryptographic secrets from the amount of power used by a device, since power consumption depends on the amount of data and the number of operations [hettwer2020applications].
@@ 3 Related Works
Adversarial training is a defense strategy aiming to train the model on both benign and adversarial samples to increase generalization capability of the model enabling it to correctly predict both benign and adversarial inputs. However, this method often comes with a trade-off, where the model sacrifices it accuracy on the benign data to increase its robustness against adversarial inputs [bai2021recent, zhao2024adversarial, zhang2025catil, jiang2024prior].
Adversarial training was first introduced in [goodfellow2014explaining], which increased the model’s robustness against FGSM attacks by sacrificing the clean accuracy of the model. Methods such as TRadeoff-inspired Adversarial DEfense via Surrogate-loss minimization (TRADES) [zhang2019theoretically] and Fair Robust Learning (FRL) [xu2021robust] have been introduced to control natural robustness-accuracy trade-off in adversarial training. Additionally, methods such as Fast Free Adversarial Training (FFAT) [shafahi2019adversarial] tackles the computational overhead of adversarial sample generation, while having a lower accuracy and robustness in comparison to standard adversarial training.
@@
Ensemble distillation-based defenses aim to harden the model by training it on soft labels and/or noisy logits instead of hard labeling, producing results via ensembles. Research works in this category include Distillation Hardening for Random Forest (DDH-RF) [apruzzese2020hardening] and SELENA [tang2022mitigating]. Hierarchical defenses, on the other hand, aim to build ensemble of models in a hierarchical manner to detect and mitigate adversarial attacks in multiple levels, which make the defense systems robust against a wider range of attacks. Hierarchical defenses include Application Constraints (AppCon) [apruzzese2020appcon], Model Stacking [salem2018ml], and Ensemble Methods as a Defense (EMD) [strauss2017ensemble].
[puigcerver2022adversarial, pavlitska2024towards] investigate the adversarial robustness of MoEs. In [han2024enhancing, zhang2025optimizing], some defense systems are proposed, which aim to protect the architecture of MoEs. Defenses such as adversarial mixture-of-expert (ADVMoE) [zhang2023robust] and synergy-of-experts (SoE) [cui2022synergy] make use of the ensemble architecture of MoEs to defend against adversarial threats. Both ADVMoE and SoE use adversarial training techniques, while training their MoE to increase the robustness of their model. Table 1 presents the differences between the category of MoE used in our defense system and other competitors.
@@ 4 Methodology
Our defense uses soft and flat structure for MoE with fractional activation of experts. In the first phase of training, we independently train a benign expert, four FGSM experts, and four PGD experts. After this pretraining stage, we move to end-to-end learning, in which the entire framework is jointly optimized. During this stage, for each batch of data, we generate adversarial variants of the inputs using both FGSM and PGD attacks, and augment them with the original benign inputs to form an expanded batch. This mixed batch preserves the diversity of benign and adversarial samples, enhancing the network’s robustness and generalization capability. Finally, the combined batch is fed into the network, so the whole framework learns both benign and adversarial patterns. Figure 4 shows an overview of our defense model’s workflow.

### v2

# Divided We Fall: Defending Against Adversarial Attacks via Soft-Gated Fractional Mixture-of-Experts with Randomized Adversarial Training

@@ Abstract
Machine learning is a powerful tool enabling full automation of a huge number of tasks without explicit programming. Despite recent progress of machine learning in different domains, these models have shown vulnerabilities when they are exposed to adversarial threats. Adversarial threats aim to hinder the machine learning models from satisfying their objectives. They can create adversarial perturbations, which are imperceptible to humans’ eyes but have the ability to cause misclassification during inference. In this paper, we propose a defense system, which devises an adversarial training module within mixture-of-experts architecture to enhance its robustness against white-box evasion attacks. In our proposed defense system, we use nine pre-trained classifiers (experts) with ResNet-18 as their backbone. During end-to-end training, the parameters of all experts and the gating mechanism are jointly updated allowing further optimization of the experts. Our proposed defense system outperforms state-of-the-art MoE-based defenses under strong white-box FGSM and PGD evaluation on CIFAR-10 and SVHN.
@@ 1 Introduction
Machine learning (ML) is a cornerstone of artificial intelligence (AI), which has been widely used in numerous applications. In computer vision, it enables face recognition and video surveillance [26, 27]. In health care, it facilitates disease diagnosis, drug repurposing, and personalized treatments [17, 44]. In natural language processing (NLP), ML enhances the quality of language translation [31, 39]. ML also empowers Large Language Models (LLMs) for code-related tasks such as code generation and understanding [22]. In cybersecurity, ML eases threat detection and manages system’s vulnerabilities [3, 56, 9]. Moreover, ML also plays an important role in other domains such as finance, speech processing, recommendation systems, autonomous vehicles, among others. However, those wide range of applications of machine learning has introduced opportunities for adversaries to exploit natural vulnerabilities of these models and launch adversarial attacks to threaten the integrity, privacy, and availability of these models [41, 8, 29, 25].
Adversarial machine learning (AML) is at intersection of cybersecurity and machine learning and aims to investigate attacks on machine learning models, and design defense systems to counter these attacks. The adversary may have some knowledge about the victim’s model and/or its training data, which facilitates the adversarial sample generation [41, 8, 29]. Figure 1 shows how an adversarial sample looks like on an MNIST [32] sample when the adversary has a perfect knowledge about the model’s parameters. Samples like this can cause model’s misclassification with a high confidence during test time.
@@
The rivalry between governments and industrial corporations has given rise to adversarial attacks causing global concerns about adversarial machine learning. For instance, the government of United States of America has established Bipartisan AI task force, which are responsible for appropriate use of AI. They introduced AI guidelines such as accuracy, reliability, robustness, privacy, transparency, and safety. They also labeled generative and adversarial AI as a threat to privacy and confidentiality [7]. Moreover, the Department of Homeland Security (DHS) has considered evasion attacks and generative AI as a national threat violating security and privacy of individuals [30]. These insights highlight the need for secure AI systems and robust ML models.
Trustworthy AI is an important domain of AI ensuring that AI models remain lawful, ethical, robust, and efficient. Robust machine learning is an important branch in trustworthy AI which aims to make the machine learning models robust against adversarial threats while remain functional in non-adversarial environments [25]. Moreover, the main focus of Robust ML is to develop defense systems that detect and/or mitigate adversarial threats while preserving their efficiency and performance to a certain extent. These defense systems could either reside in the models or serve as an external entity protecting ML models [41, 8, 29, 25].
Although adversarial machine learning encompasses multiple threat categories, including poisoning, model extraction, and privacy attacks, this study focuses specifically on inference-time robustness against white-box evasion attacks. Our evaluation therefore targets gradient-based adversarial examples generated with full model access. Training-time attacks and broader privacy or extraction threats are outside the scope of this work and are left for future investigation.
In pursuit of building a secure system, we build a novel defense system named Divided We Fall (DWF). In DWF, we employ Mixture of Experts (MoE) architecture which consists of benign and adversarial experts. The training scheme in DWF is joint, meaning that we do not freeze the weights of our pretrained experts during end-to-end training, which leads to stronger generalization capability of our proposed defense system improving the clean accuracy and robustness of our model. To ensure rigorous evaluation, robustness is assessed using strong white-box attacks, including properly tuned multi-step PGD with many iterations, multi-step restarts, and adaptive variants that optimize through the full mixture-of-experts architecture.
The proposed method is termed Divided We Fall to emphasize its distributed robustness principle. Rather than relying on a single monolithic classifier, which can be entirely compromised by adversarial perturbations, DWF decomposes the model into multiple specialized experts trained on different regimes (benign, FGSM, and PGD). The soft gating mechanism adaptively combines their outputs at inference time, allowing the system to remain reliable even if individual experts are partially affected. By avoiding a single point of failure and encouraging specialization and cooperation, the overall model becomes more resilient to adversarial attacks.
@@ 2 Background and Problem Setting
In this section, we introduce MoE architecture and a background on adversarial threat model.
2.1 Mixture-of-Experts
MoE is an ensemble architecture using multiple experts, a gating mechanism to assign weights, and a router to aggregate results [37, 10, 43]. MoEs can be categorized based on their gating mechanism, routing topology, training strategy, and experts activation, which is shown in Figure 2.
@@ 2.1 Mixture-of-Experts
Based on gating mechanism, MoE could be either soft or hard; in soft MoE the gating assigns non-zero weights to each expert forming fractional activation of experts, while in hard MoE, the gating mechanism chooses a subset of experts and assign wights to them making experts activation as sparse, or single if only one of them is chosen [37, 10, 43].
Training strategy could be either joint or frozen. In the first one, both router and experts are trained simultaneously, while the latter one uses pretrained experts and freezes their weights during training of the router [37, 10, 43].
Based on the routing topology we have flat and hierarchical routing schemes. In the flat MoE, all the experts reside in a single layer and the router chooses from them. In the hierarchical topology, the first router groups the experts into subgroups, and, then, the second one chooses some experts from these subgroups [37, 10, 6].
Using an expert-driven model allows for specialization within each part of the system, much like Unmanned Aerial Vehicle (UAVs) assigned specific roles within a swarm. These experts could respond dynamically to different types of attacks or threats, providing targeted defenses that are optimized for specific conditions, rather than relying on a single, monolithic defense mechanism [6].
2.2 General Adversarial Threat Models
Adversarial threats can be viewed from different angles: attacker’s knowledge, type, specificity, and objective.
2.2.1 Attacker’s Knowledge
Attacker’s knowledge shows the access level of the adversary to model’s parameters, gradients, architecture, and training data. This knowledge is categorized into white-box, black-box, and gray-box based on the adversary’s access level [41, 8].
White Box
Adversaries with white-box knowledge are the most effective and destructive attackers. These attackers have a full access to model’s architecture, parameters, and gradients during training with full knowledge on the training data; this enables the attackers to generate adversarial samples based on the gradients and training data, which makes them very powerful. However, in a real-world scenario, adversaries can’t achieve such an access, making them more suitable for research and testing the models’ robustness [41, 8].
Black Box
In black-box scenario, unlike white-box, the adversaries have no knowledge about the model and the presence of any defense system. The adversaries can only give inputs and receive the corresponding outputs [41, 8, 33].
Gray Box
Gray box knowledge lies between white box knowledge and black box knowledge. The adversary has some degree of knowledge about the target system, which is not as perfect as white box, and is not as scarce as black box knowledge [41, 8].
2.2.2 Attacker’s specificity
Based on specificity, each attack could either be targeted or untargeted. In targeted attacks, the adversary aims to degrade the model’s performance when it faces specific patterns/inputs, while in untargeted attacks, also named to as indiscriminate, the goal is to cause any misclassification [41, 8].
2.2.3 Attacker’s objective
Based on the objective, the adversary can follow degrade, breach, or both scenarios. In the first scenario, the adversary aims to harm the system’s performance and/or violets its integrity. In breach, the adversary aims to explore the model in order to steal sensitive information, without affecting the model’s performance [41, 8].
2.2.4 Attacker’s type
@@
Evasion
An evasion attack targets the model at test time without affecting the system’s integrity on clean data. Evasion attacks are the most practical attack types aiming to degrade the model’s performance by causing misclassification during testing [41, 8, 15]. For instance attacks such as Fast Gradient Sign Method (FGSM) [16], Projected Gradient Descent (PGD) [36], and Carlini & Wagner (C&W) [11] make use of the model’s gradient during training to generate samples which evade the classifier during the test time. Equations 3 and 4, in Section 5, show how an adversarial sample is generated in FGSM and PGD, respectively.
Other Attack Types
Poisoning attacks compromise a model’s integrity by corrupting training data or embeddings, either through label manipulation or clean label attacks [41, 8, 55, 5, 59, 65]. Exploratory attacks aim to violate privacy of the system. These attacks probe the target model to gain information about model’s parameters, training data, and algorithms on without affecting the performance and integrity of the model [46, 62, 50]. Moreover, physical attacks are launched on the models, which are deployed on a special physical hardware such as FPGA, GPU, and edge devices to exploit vulnerabilities emerged due to the increased attack surface in the physical layer [45, 20]. In this research work, we only focus on defending against evasion attacks to assess the resilience of our proposed approach under severe white-box attacks.
Figure 3 shows an overview of adversarial threat model.
@@ 3 Related Works
Adversarial training is a defense strategy aiming to train the model on both benign and adversarial samples to increase generalization capability of the model enabling it to correctly predict both benign and adversarial inputs. However, this method often comes with a trade-off, where the model sacrifices it accuracy on the benign data to increase its robustness against adversarial inputs [4, 64, 63, 23].
Adversarial training was first introduced in [16], which increased the model’s robustness against FGSM attacks by sacrificing the clean accuracy of the model. Methods such as TRadeoff-inspired Adversarial DEfense via Surrogate-loss minimization (TRADES) [58] and Fair Robust Learning (FRL) [57] have been introduced to control natural robustness-accuracy trade-off in adversarial training. Additionally, methods such as Fast Free Adversarial Training (FFAT) [49] tackles the computational overhead of adversarial sample generation, while having a lower accuracy and robustness in comparison to standard adversarial training.
@@
Ensemble distillation-based defenses aim to harden the model by training it on soft labels and/or noisy logits instead of hard labeling, producing results via ensembles. Research works in this category include Distillation Hardening for Random Forest (DDH-RF) [1] and SELENA [54]. Hierarchical defenses, on the other hand, aim to build ensemble of models in a hierarchical manner to detect and mitigate adversarial attacks in multiple levels, which make the defense systems robust against a wider range of attacks. Hierarchical defenses include Application Constraints (AppCon) [2], Model Stacking [47], and Ensemble Methods as a Defense (EMD) [52].
[42, 40] investigate the adversarial robustness of MoEs. In [18, 60], some defense systems are proposed, which aim to protect the architecture of MoEs. Defenses such as adversarial mixture-of-expert (ADVMoE) [61] and synergy-of-experts (SoE) [13] make use of the ensemble architecture of MoEs to defend against adversarial threats. Both ADVMoE and SoE use adversarial training techniques, while training their MoE to increase the robustness of their model.
Table 1 presents the differences between the category of MoE used in our defense system and other competitors.
@@
While standard adversarial training methods improve robustness by augmenting training data with adversarial examples, they typically train a single model with a uniform objective. In contrast, our approach explicitly decomposes the problem into multiple specialized experts, each trained on different regimes (benign, FGSM, and PGD), and combines them via a soft gating mechanism. This distinguishes totally from conventional adversarial training, which lacks explicit expert specialization and input-dependent routing.
@@ 4 Methodology
Our defense uses soft and flat structure for MoE with fractional activation of experts. In the first phase of training, we independently train a classifier on benign data (benign expert), four classifiers on FGSM data (FGSM experts), and four classifiers on PGD data (PGD experts). This design encourages diversity among experts, which is a key principle in mixture-of-experts architectures. After this pretraining stage, we move to end-to-end learning, in which the entire framework is jointly optimized. During this stage, for each batch of data, we generate adversarial variants of the inputs using both FGSM and PGD attacks, and augment them with the original benign inputs to form an expanded batch. This mixed batch preserves the diversity of benign and adversarial samples, enhancing the network’s robustness and generalization capability. Finally, the combined batch is fed into the network, so the whole framework learns both benign and adversarial patterns. Figure 4 shows an overview of our defense model’s workflow.
@@
During training, we sample attack strengths (FGSM $\epsilon$ and PGD iterations) from a range rather than fixing a single configuration. This strategy prevents overfitting to a specific perturbation magnitude and exposes the model to adversarial examples of varying difficulty, improving robustness generalization across threat levels. Consequently, the model is not tuned to a particular evaluation $\epsilon$ but learns to remain stable under a spectrum of perturbations. A detailed ablation over all sampling ranges is left for future investigation due to computational cost.
@@
The gating function is implemented as a lightweight learnable router that takes the same input $x$ as the experts and produces a soft assignment vector using a small multilayer perceptron followed by a softmax normalization. The final prediction is obtained via a weighted sum of all expert logits (soft routing), rather than hard or sparse selection. All experts and the gating network are trained jointly in an end-to-end manner using the standard cross-entropy loss. No auxiliary routing objectives, load-balancing losses, temperature scaling, or sparsity regularization are used. This simple design avoids additional hyperparameters and improves reproducibility.
The weights $w_{i}(x)$ in (1) are generated by a learnable gating network $g(.)$. Given an input $x$, the gating network produces a score vector $s(x)\in\mathbb{R}^{n}$, which is normalized using a softmax function to obtain non-negative weights that sum to one:
$w_{i}(x)=\frac{e^{s_{i}(x)}}{\sum_{j=1}^{n}e^{s_{j}(x)}}$ (2)
The gating network is trained jointly with all experts during end-to-end training by backpropagating the cross-entropy loss through both the experts and the gating mechanism. Thus, $w(x)$ is not fixed or heuristic, but learned from data in an input-dependent manner.
During training, we randomize attack parameters such as perturbation magnitude and step size. This randomness is not strictly required for robustness, but is used to increase perturbation diversity and avoid overfitting to a single fixed attack configuration. By exposing the experts to a wider range of adversarial strengths, the model generalizes better to unseen attacks. In practice, similar behavior can still be achieved with fixed parameters, although randomized training typically provides slightly more stable robustness.
The number of experts determines a trade-off between robustness and computational cost. A larger number of experts enables greater specialization to different input regimes, which can improve robustness. However, both memory and inference cost grow approximately linearly with the number of experts, while the marginal gain diminishes as experts begin to learn overlapping representations. Therefore, we adopt nine experts as a practical compromise that provides strong robustness while maintaining reasonable computational efficiency.
All experts share the same backbone architecture (ResNet-18) in our implementation. This choice is intentional to ensure fair comparisons and to isolate the contribution of the proposed soft-gated mixture-of-experts training strategy, rather than improvements arising from heterogeneous model capacities. Despite identical architectures, diversity is still achieved through specialization: each expert is trained on different input regimes (benign and adversarial attacks), which results in distinct decision boundaries and complementary behaviors. Incorporating heterogeneous backbones could further increase diversity and potentially improve robustness; however, this would introduce additional computational cost and confounding factors. Exploring architectural diversity among experts is an interesting direction for future work.